# EDEN: Exploring Disks Embedded in N-body simulations of Milky-Way-mass halos from Symphony

### v1

@@ Abstract
facilities: Stanford Research Computing Center, SLAC Data Facility, TACC Stampede2software: astropy (Astropy Collaboration et al., 2013, 2018), matplotlib, arXiv, ADS
@@ 1 Introduction
The enhanced mass loss effect from the host galaxy is primarily studied using pairs of hydrodynamic simulations and DMO counterparts starting out from the same initial conditions, focusing on satellites around MW-mass halos (Sawala et al., 2013; Brooks & Zolotov, 2014; Sawala et al., 2015; Wetzel et al., 2016; Zhu et al., 2016; Despali & Vegetti, 2017; Sawala et al., 2017; Garrison-Kimmel et al., 2017; Graus et al., 2018; Samuel et al., 2020; Barry et al., 2023; Jung et al., 2024). It is revealed that significant suppression of satellite and subhalo abundances occur, especially near host centers for subhalos with smaller pericenters and more radially-biased velocities due to the addition of baryons (stars and gas) compared to the DMO case. Since this effect is primarily due to the central galaxy’s mass concentration, it is, in principle, agnostic towards the specific treatment of star formation and feedback physics in the hydrodynamic simulations.
@@
To further isolate the pure gravitational effect of the central galaxy, several simulation works further combined analytic disk potentials with DMO zoom-in simulations (D’Onghia et al., 2010; Errani et al., 2017; Garrison-Kimmel et al., 2017; Kelley et al., 2019). These explorations also discovered similar levels of subhalo abundance suppression as hydrodynamic simulations when compared to their DMO counterparts. These studies interpreted reduced subhalo abundances as complete physical disruption and thus concluded that the central galaxy alone could potentially solve the missing satellites and TBTF problems. [^1] The latter problem is addressed because heavier subhalos experience stronger dynamical friction and sink more quickly toward halo centers, where the disk effects are strongest. However, we will argue that reduced subhalo abundances above a peak (or present-day) mass threshold is not evidence of physical disruption; rather, it indicates enhanced mass loss which is in agreement with recent semi-analytic studies (Green et al., 2022). Furthermore, previous studies often implicitly assume that satellite galaxies hosted by subhalos would “disrupt” along with the subhalo itself, which is also an oversimplification as “orphan” galaxies out-surviving their subhalos are required to match observations (Behroozi et al., 2019) and depends on the specific resolution of the simulation used (Mansfield et al., 2023). Thus, we are motivated to revisit how disk potentials affect subhalo populations in both the MW and “typical” MW-mass halos.
@@
The MW lives in a halo that has a virial mass of $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ (Bland-Hawthorn & Gerhard, 2016; McMillan, 2017; Gaia Collaboration et al., 2018; Posti & Helmi, 2019; Watkins et al., 2019; Cautun et al., 2020; Labini et al., 2023). All the DMO-embedded disk simulations mentioned above focused on MW-mass halos with $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ and embedded massive disk potentials similar to our Galaxy ($M_{\rm d}\gtrsim 5\times 10^{10}\mathrm{M_{\astrosun}}$, Licquia & Newman, 2015; Bland-Hawthorn & Gerhard, 2016; McMillan, 2017; Cautun et al., 2020). Furthermore, in simulations such as D’Onghia et al. (2010) (fixed disk-to-halo mass ratio) and Phat ELVIS (abundance matching $M_{\ast}$ to $M_{\rm vir}$, Behroozi et al., 2013a), the simplistic disk growth models adopted does not account for diversity in halo assembly and stellar mass growth in a correlated manner. Although Garrison-Kimmel et al. (2017) used more physical disk formation histories from their hydrodynamic runs, their sample size was limited to only two MW-mass hosts. Therefore, a large parameter space of disk-to-halo mass ratios and disk growth histories is yet to be explored for MW-mass halos in terms of their subhalo abundance suppression effects.
@@
In this work, we aim to fill this gap by better quantifying subhalo abundances over a wider range of disk masses and disk formation histories in MW-mass halos. We present the EDEN (Exploring Disks Embedded in N-body) simulations where we re-simulate 45 $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ MW-mass halos from the Symphony (Nadler et al., 2023b) zoom-in simulation compilation with embedded analytic disk potentials. Two technical highlights of this work are that: i) EDEN evolves disk growth according to the empirical galaxy–halo connection model UniverseMachine (UM hereafter; Behroozi et al., 2019; Wang et al., 2021) that explicitly correlates the star formation history with the halo assembly history; ii) EDEN uses the state-of-the-art particle-tracking subhalo finder Symfind (Mansfield et al., 2023) to enable more robust subhalo tracking. We investigate the reduction in the abundance of subhalos as a function of host mass, subhalo mass, 3D and pericenter distances to their hosts, and infall times. As the disk mass has a dominant impact on subhalo mass loss rates over other factors like disk radius, scale height, and density profiles (Garrison-Kimmel et al., 2017; Green et al., 2022), we further re-simulate nine of our 45 zoom-in halos with ×$\times 2.5$ heavier disk potentials while keeping the form of their disk growth histories fixed. This subset of higher-disk-mass simulations, along with the main set of EDEN simulations, provides good sampling of subhalo abundance suppression over the full range of central galaxy stellar masses for typical $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ MW-mass halos.
@@ 2 Methodology
Here, we present the methodology for embedding disk potentials into the SymphonyMilkyWay DMO simulations. Section 2.1 summarizes basic properties of the Symphony compilation (Nadler et al., 2023b); Section 2.2 introduces how we apply the galaxy–halo model UM to Symphony and retrieve the stellar mass growth histories of the MW central galaxies. Section 2.3 summarizes the disk density components and how we embed them into EDEN host halos at $z=3$; in Section 2.4 introduces the nine-host subsample that is re-simulated with 2.5×2.5$\times$2.5 × higher disk mass.
@@ 2.1 SymphonyMilkyWay dark-matter-only zoom-ins
In this work, we focus on the 45 MW-mass host halos from Symphony. Originally, these halos were selected from the Chinchilla c125-2048 cosmological simulation (Mao et al., 2015) with masses $M_{\rm vir}/\mathrm{M_{\astrosun}}\in 10^{12.1\pm 0.03}$. Resolution is increased within ‘zoomed-in’ Lagrangian regions that contain particles within $\sim 10R_{\rm vir}$ of the MW hosts at $z=0$. The initial conditions for the zoom-in regions are generated with the public code Music (Hahn & Abel, 2011). [^2] https://www-n.oca.eu/ohahn/MUSIC/ Halos were then re-simulated from $z=99$ using Gadget-2 (Springel, 2005) [^3] https://wwwmpa.mpa-garching.mpg.de/gadget/. All the simulations in this suite have particle masses of $m_{\rm DM}=4\times 10^{5}\mathrm{M_{\astrosun}}$ and adopt a comoving Plummer-equivalent force softening scale of $\epsilon=243$ pc in the high-resolution zoom-in region. This mass resolution is equivalent to having $8192^{3}$ particles in the 125 $h^{-1}$pc parent box c125-2048. In EDEN, we re-simulate these systems with the same particle resolution and force softening while embedding disk potentials (see Section 2.3).
Halos in Symphony were identified using the phase-space halo finder Rockstar (Behroozi et al., 2013b). [^4] https://bitbucket.org/gfcstanford/rockstar/src/main/ We use the Bryan & Norman (1998) virial overdensity definition, which is equivalent to an overdensity of $\Delta_{c}=99.2$ at $z=0$. Halo merger trees are constructed using consistent-trees (Behroozi et al., 2013c). [^5] https://bitbucket.org/pbehroozi/consistent-trees/src/main/ We use these consistent-trees catalogs to track subhalos before they cross into their hosts’ virial radii ($R_{\rm vir,host}$). We use the Symfind (Mansfield et al., 2023) particle-tracking subhalo finder after subhalo infall for more reliable subhalo tracking, which is a major improvement of this work over previous non-particle-tracking halo finders used to analyze embedded-disk simulations (e.g., AHF,  Knollmann & Knebe 2009, for Garrison-Kimmel et al. 2017 and Rockstar for Kelley et al. 2019; Nadler et al. 2023b). All halo catalogs and merger trees for Symphony, including Symfind subhalo catalogs, are publicly available at https://web.stanford.edu/group/gfc/symphony/. The halo catalogs of the EDEN simulations presented in this work will be made public on the same website upon publication of this manuscript.
@@ 2.2 UniverseMachine star formation histories
In the left panel of Fig.1, we show the MW host halo masses ($M_{\rm vir,host}$), the UM-predicted stellar (disk) masses for the MW galaxy ($M_{\ast,\rm Disk}$), the stellar half-mass formation scale factor ($a_{\ast,1/2}$), and the halo half-mass formation scale factor ($a_{\rm vir,1/2}$). We assume that the three-component disk potential (Section 2.3) contains the total stellar mass predicted by UM, and we use $M_{\ast}$ and $M_{\ast,\rm Disk}$ interchangeably. Compared to the latest observational constraints on the MW’s halo mass (Bland-Hawthorn & Gerhard, 2016; McMillan, 2017; Gaia Collaboration et al., 2018; Posti & Helmi, 2019; Watkins et al., 2019; Cautun et al., 2020; Labini et al., 2023), the SymphonyMilkyWay hosts have quite representative virial masses (mean $M_{\rm vir}\sim 10^{12.1}\mathrm{M_{\astrosun}}$). However, the MW has an atypically large ($\gtrsim 1\sigma$) stellar mass (Licquia & Newman, 2015; Bland-Hawthorn & Gerhard, 2016; McMillan, 2017; Cautun et al., 2020) for a halo of $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$, and is larger than the average empirical SMHM relation from UM (Wang et al., 2021). We also show observational measurements of the halo mass (Watkins et al., 2010) and stellar mass (Tamm et al., 2012) of M31, which is even higher stellar-to-halo mass ratio. Previous embedded disk DMO simulations marked in orange from Garrison-Kimmel et al. (2017) and Kelley et al. (2019) all focused on MW-like high-disk-mass systems ($\gtrsim 1\sigma$ up scatter) for MW-mass halos around $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$. Therefore, a large part of the $M_{\ast,\rm Disk}/M_{\rm vir,host}$ parameter space that the majority of MW-mass galaxies find themselves in has not yet been explored. EDEN aims to fill this important gap and provide a more comprehensive understanding of disk effects on subhalo abundances in MW-mass hosts with a range of galaxy formation histories.
Besides covering a much broader range of disk-to-halo mass ratios than previous literature, EDEN also accounts for the intrinsic scatter in disk formation histories due to the diversity in halo assembly. This is naturally achieved with the UM-predicted stellar mass growth histories of the disks, which includes a correlation between galaxy and halo growth. In Fig. 1, we also show the wide distribution of star formation half mass scale ($a_{\ast,1/2}$) ranging from 0.4 to 0.9, which anti-correlates with stellar mass and halo concentration $c_{\rm vir}$ (middle panel), and are mostly later than their halo half-mass formation time scale (right panel). Disks with larger stellar masses also form earlier, giving them more time to tidally strip subhalos historically.
@@ 2.3 Embedded disk potential setup
Here, $G$ is the gravitational constant, $M_{d}$ is the disk mass for stellar ($0.60M_{\ast}$) or gaseous ($0.27M_{\ast}$) disk, $R_{d}$ is the disk scale radius (2.5 kpc for stellar and 7 kpc for gaseous at $z=0$), and $h_{d}$ is the disk scale height (0.35 kpc for stellar disk and 0.084 kpc for gaseous at $z=0$). The gas fraction is consistent with the latest empirical constraints on the HI to stellar mass ratio (Padmanabhan & Loeb, 2020), which yields a gas-to-mass ratio of $\sim 30\%$ at $z>2$ and $\sim 15\%$ at $z<1$ for MW-mass halos (Fig. 1 in Padmanabhan & Loeb 2020). Our stellar disk scale radius is consistent with the MW’s thin ($2.0\pm 0.2$ kpc) and thick ($2.6\pm 0.5$ kpc) stellar disk constraints (Bland-Hawthorn & Gerhard, 2016). The MW gaseous disk is less constrained with typically 2$\sim$ 2$\times$ 2 × the radius of the stellar disk (Bland-Hawthorn & Gerhard, 2016; Posti & Helmi, 2019), indicating our assumption is reasonable. Our choice of the stellar disk scale height is representative of MW measurements that span from 140 to 430 pc using different stellar populations (McKee et al., 2015), while the gaseous disk scale height assumed is consistent with cold CO disk measurements of $\lesssim 100$ pc (Heyer & Dame, 2015). In addition to the CO disk, there is also a thin and a thick cold HI disk, both having radially varying scale heights ($h>200$ pc, Valenti et al. 2016), and we do not model this subtlety in this work.
For the spherical bulge ($\Phi_{\rm b}$), since it is only $\sim 10\%$ of the total disk mass, its specific density profile would not change our results significantly. Hence, without loss of generality, we assume a Hernquist potential profile (Hernquist, 1990) for it:

### v3

@@ Abstract
facilities: Stanford Research Computing Center, SLAC Data Facility, TACC Stampede2software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), arXiv, ADS
@@ 1 Introduction
The enhanced mass loss effect from the host galaxy is primarily studied using pairs of hydrodynamic simulations and DMO counterparts starting out from the same initial conditions, focusing on satellites around MW-mass halos (Sawala et al., 2013; Brooks & Zolotov, 2014; Sawala et al., 2015; Wetzel et al., 2016; Zhu et al., 2016; Despali & Vegetti, 2017; Sawala et al., 2017; Garrison-Kimmel et al., 2017; Graus et al., 2018; Samuel et al., 2020; Barry et al., 2023; Jung et al., 2024). It is revealed that significant suppression of satellite and subhalo abundances occurs, especially near host centers for subhalos with smaller pericenters and more radially-biased velocities due to the addition of baryons (stars and gas) compared to the DMO case. Since this effect is primarily due to the central galaxy’s mass concentration, it is, in principle, agnostic towards the specific treatment of star formation and feedback physics in the hydrodynamic simulations.
@@
To further isolate the pure gravitational effect of the central galaxy, several simulation works further combined analytic disk potentials with DMO zoom-in simulations (D’Onghia et al., 2010; Errani et al., 2017; Garrison-Kimmel et al., 2017; Kelley et al., 2019). These explorations also discovered similar levels of subhalo abundance suppression as hydrodynamic simulations when compared to their DMO counterparts. These studies interpreted reduced subhalo abundances as complete physical disruption and thus concluded that the central galaxy alone could potentially solve the missing satellites and TBTF problems. [^1] The latter problem is addressed because heavier subhalos experience stronger dynamical friction and sink more quickly toward halo centers, where the disk effects are strongest. However, we will argue that reduced subhalo abundances above a peak (or present-day) mass threshold is not evidence of physical disruption; rather, it indicates enhanced mass loss which is in agreement with recent semi-analytic studies (Green et al., 2022). Furthermore, previous studies often implicitly assume that satellite galaxies hosted by subhalos would “disrupt” along with the subhalo itself, which is also an oversimplification as “orphan” galaxies out-surviving their subhalos are required to match observations (Behroozi et al., 2019) and depend on the specific resolution of the simulation used (Mansfield et al., 2024). Thus, we are motivated to revisit how disk potentials affect subhalo populations in both the MW and “typical” MW-mass halos.
@@
The MW lives in a halo that has a virial mass of $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ (Bland-Hawthorn & Gerhard, 2016; McMillan, 2017; Gaia Collaboration et al., 2018; Posti & Helmi, 2019; Watkins et al., 2019; Cautun et al., 2020; Labini et al., 2023). All the DMO-embedded disk simulations mentioned above focused on MW-mass halos with $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ and embedded massive disk potentials similar to our Galaxy ($M_{\rm d}\gtrsim 5\times 10^{10}\mathrm{M_{\astrosun}}$, Licquia & Newman, 2015; Bland-Hawthorn & Gerhard, 2016; McMillan, 2017; Cautun et al., 2020). Furthermore, in simulations such as D’Onghia et al. (2010) (fixed disk-to-halo mass ratio) and Phat ELVIS (abundance matching $M_{\ast}$ to $M_{\rm vir}$, Behroozi et al., 2013a), the simplistic disk growth models adopted do not account for diversity in halo assembly and stellar mass growth in a correlated manner. Although Garrison-Kimmel et al. (2017) used more physical disk formation histories from their hydrodynamic runs, their sample size was limited to only two MW-mass hosts. Therefore, a large parameter space of disk-to-halo mass ratios and disk growth histories is yet to be explored for MW-mass halos in terms of their subhalo abundance suppression effects.
@@
In this work, we aim to fill this gap by better quantifying subhalo abundances over a wider range of disk masses and disk formation histories in MW-mass halos. We present the EDEN (Exploring Disks Embedded in N-body) simulations where we re-simulate 45 $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ MW-mass halos from the Symphony (Nadler et al., 2023b) zoom-in simulation compilation with embedded analytic disk potentials. Two technical highlights of this work are that: i) EDEN evolves disk growth according to the empirical galaxy–halo connection model UniverseMachine (UM hereafter; Behroozi et al., 2019; Wang et al., 2021) that explicitly correlates the star formation history with the halo assembly history; ii) EDEN uses the state-of-the-art particle-tracking subhalo finder Symfind (Mansfield et al., 2024) to enable more robust subhalo tracking. We investigate the reduction in the abundance of subhalos as a function of host mass, subhalo mass, 3D and pericenter distances to their hosts, and infall times. As the disk mass has a dominant impact on subhalo mass loss rates over other factors like disk radius, scale height, and density profiles (Garrison-Kimmel et al., 2017; Green et al., 2022), we further re-simulate nine of our 45 zoom-in halos with ×$\times 2.5$ heavier disk potentials while keeping the form of their disk growth histories fixed. This subset of higher-disk-mass simulations, along with the main set of EDEN simulations, provides good sampling of subhalo abundance suppression over the full range of central galaxy stellar masses for typical $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ MW-mass halos.
@@ 2 Methodology
Here, we present the methodology for embedding disk potentials into the SymphonyMilkyWay DMO simulations. Section 2.1 summarizes basic properties of the Symphony compilation (Nadler et al., 2023b); Section 2.2 introduces how we apply the galaxy–halo model UM to Symphony and retrieve the stellar mass growth histories of the MW central galaxies. Section 2.3 summarizes the disk density components and how we embed them into EDEN host halos at $z=3$; in Section 2.4 we introduce the nine-host subsample that is re-simulated with 2.5×2.5$\times$2.5 × higher disk mass.
@@ 2.1 SymphonyMilkyWay dark-matter-only zoom-ins
In this work, we focus on the 45 MW-mass host halos from Symphony. Originally, these halos were selected from the Chinchilla c125-2048 cosmological simulation (Mao et al., 2015) with masses $M_{\rm vir}/\mathrm{M_{\astrosun}}\in 10^{12.1\pm 0.03}$. Resolution is increased within ‘zoomed-in’ Lagrangian regions that contain particles within $\sim 10R_{\rm vir}$ of the MW hosts at $z=0$. The initial conditions for the zoom-in regions are generated with the public code Music (Hahn & Abel, 2011). [^2] https://www-n.oca.eu/ohahn/MUSIC/ Halos were then re-simulated from $z=99$ using Gadget-2 (Springel, 2005) [^3] https://wwwmpa.mpa-garching.mpg.de/gadget/. All the simulations in this suite have particle masses of $m_{\rm DM}=4\times 10^{5}\mathrm{M_{\astrosun}}$ and adopt a comoving Plummer-equivalent force softening scale of $\epsilon=243$ pc in the high-resolution zoom-in region. This mass resolution is equivalent to having $8192^{3}$ particles in the 125 $h^{-1}$ Mpc parent box c125-2048. In EDEN, we re-simulate these systems with the same particle resolution and force softening while embedding disk potentials (see Section 2.3).
Halos in Symphony were identified using the phase-space halo finder Rockstar (Behroozi et al., 2013b). [^4] https://bitbucket.org/gfcstanford/rockstar/src/main/ We use the Bryan & Norman (1998) virial overdensity definition, which is equivalent to an overdensity of $\Delta_{c}=99.2$ at $z=0$. Halo merger trees are constructed using consistent-trees (Behroozi et al., 2013c). [^5] https://bitbucket.org/pbehroozi/consistent-trees/src/main/ We use these consistent-trees catalogs to track subhalos before they cross into their hosts’ virial radii ($R_{\rm vir,host}$). We use the Symfind (Mansfield et al., 2024) particle-tracking subhalo finder after subhalo infall for more reliable subhalo tracking, which is a major improvement of this work over previous non-particle-tracking halo finders used to analyze embedded-disk simulations (e.g., AHF,  Knollmann & Knebe 2009, for Garrison-Kimmel et al. 2017 and Rockstar for Kelley et al. 2019; Nadler et al. 2023b). All halo catalogs and merger trees for Symphony, including Symfind subhalo catalogs, are publicly available at https://web.stanford.edu/group/gfc/symphony/. The EDEN halo catalogs (https://web.stanford.edu/group/gfc/gfcsims/build/html/eden_overview.html) can be accessed as part of the Symphony data products and its simulation code repository (https://github.com/Bulk826R/EDEN) is publicly available for download.
@@ 2.2 UniverseMachine star formation histories
In the left panel of Fig.1, we show the MW host halo masses ($M_{\rm vir,host}$), the UM-predicted stellar (disk) masses for the MW galaxy ($M_{\ast,\rm Disk}$), the stellar half-mass formation scale factor ($a_{\ast,1/2}$), and the halo half-mass formation scale factor ($a_{\rm vir,1/2}$). We assume that the three-component disk potential (Section 2.3) contains the total stellar mass predicted by UM, and we use $M_{\ast}$ and $M_{\ast,\rm Disk}$ interchangeably. Compared to the latest observational constraints on the MW’s halo mass (Bland-Hawthorn & Gerhard, 2016; McMillan, 2017; Gaia Collaboration et al., 2018; Posti & Helmi, 2019; Watkins et al., 2019; Cautun et al., 2020; Labini et al., 2023), the SymphonyMilkyWay hosts have quite representative virial masses (mean $M_{\rm vir}\sim 10^{12.1}\mathrm{M_{\astrosun}}$). However, the MW has an atypically large ($\gtrsim 1\sigma$) stellar mass (Licquia & Newman, 2015; Bland-Hawthorn & Gerhard, 2016; McMillan, 2017; Cautun et al., 2020) for a halo of $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$, and is larger than the average empirical SMHM relation from UM (Wang et al., 2021). We also show observational measurements of the halo mass (Watkins et al., 2010) and stellar mass (Tamm et al., 2012) of M31, which has even higher stellar-to-halo mass ratio. Previous embedded disk DMO simulations marked in orange from Garrison-Kimmel et al. (2017) and Kelley et al. (2019) all focused on MW-like high-disk-mass systems ($\gtrsim 1\sigma$ up scatter) for MW-mass halos around $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$. Therefore, a large part of the $M_{\ast,\rm Disk}/M_{\rm vir,host}$ parameter space that the majority of MW-mass galaxies find themselves in has not yet been explored. EDEN aims to fill this important gap and provide a more comprehensive understanding of disk effects on subhalo abundances in MW-mass hosts with a range of galaxy formation histories.
Besides covering a much broader range of disk-to-halo mass ratios than previous literature, EDEN also accounts for the intrinsic scatter in disk formation histories due to the diversity in halo assembly. This is naturally achieved with the UM-predicted stellar mass growth histories of the disks, which include a correlation between galaxy and halo growth. In Fig. 1, we also show the wide distribution of star formation half mass scale ($a_{\ast,1/2}$) ranging from 0.4 to 0.9, which anti-correlates with stellar mass and halo concentration $c_{\rm vir}$ (middle panel), and are mostly later than their halo half-mass formation time scale (right panel). Disks with larger stellar masses also form earlier, giving them more time to tidally strip subhalos historically.
@@ 2.3 Embedded disk potential setup
Here, $G$ is the gravitational constant, $M_{d}$ is the disk mass for ‘stellar’ ($0.60M_{\ast}$) or ‘gaseous’ ($0.27M_{\ast}$) disk, $R_{d}$ is the disk scale radius (2.5 kpc for stellar and 7 kpc for gaseous at $z=0$), and $h_{d}$ is the disk scale height (0.35 kpc for stellar disk and 0.084 kpc for gaseous at $z=0$). The gas fraction is consistent with the latest empirical constraints on the HI to stellar mass ratio (Padmanabhan & Loeb, 2020), which yields a gas-to-mass ratio of $\sim 30\%$ at $z>2$ and $\sim 15\%$ at $z<1$ for MW-mass halos (Fig. 1 in Padmanabhan & Loeb 2020). As found in Garrison-Kimmel et al. (2017), the inclusion of a gas disk potential that follows the gas disk growth in the hydrodynamic run is less important for determining subhalo abundances than the mass of the stellar disk. Motivated by this finding, we assume a fixed fraction of the UM-predicted stellar mass to be in the ‘gaseous’ disk from $z=3$ to $z=0$. We re-simulate three EDEN hosts in Appendix F with gas disks growing more realistically following the predictions of NeutralUniverseMachine (Guo et al., 2023), an empirical gas–halo connection model calibrated on UM (Behroozi et al., 2019). We find that the gas disk potential is indeed less consequential than the mass normalization of the stellar disk, in line with Garrison-Kimmel et al. (2017).
As for the disk size, the stellar disk scale radius is consistent with the MW’s thin ($2.0\pm 0.2$ kpc) and thick ($2.6\pm 0.5$ kpc) stellar disk constraints (Bland-Hawthorn & Gerhard, 2016). The MW gaseous disk is less constrained with typically 2$\sim$ 2$\times$ 2 × the radius of the stellar disk (Bland-Hawthorn & Gerhard, 2016; Posti & Helmi, 2019), indicating our assumption is reasonable. Our choice of the stellar disk scale height is representative of MW measurements that span from 140 to 430 pc using different stellar populations (McKee et al., 2015), while the gaseous disk scale height assumed is consistent with cold CO disk measurements of $\lesssim 100$ pc (Heyer & Dame, 2015). In addition to the CO disk, there is also a thin and a thick cold HI disk, both having radially varying scale heights ($h>200$ pc, Valenti et al. 2016), and we do not model this subtlety in this work.
Since the spherical bulge ($\Phi_{\rm b}$) only contains $\sim 10\%$ of the total disk mass, it is unlikely that our choice of its density profile would significantly change our results. Hence, without loss of generality, we assume a Hernquist potential for the bulge profile (Hernquist, 1990):